OpenStudy (anonymous):

13. Use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth. X^2+3= -4x

OpenStudy (anonymous):

a.1, 3 b. –1, –3 c. 1, –3 d. 1, –3

jhonyy9 (jhonyy9):

x^2 +4x +3 =0 (x+3)(x+1)=0 so here is factorized after this can you solve it easy ,yes ?

OpenStudy (anonymous):

@jhonyy9 , they said to use the quadratic formula. @bubble1860 , the solutions to \(0=ax^2+bx+c\) are given by the quadratic formula as follows.\[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

jhonyy9 (jhonyy9):

yes is right but in this way is more ,much easy ...

OpenStudy (anonymous):

The question asks to use the quadratic formula.